OpenStudy (jiteshmeghwal9):

If \(\alpha\) & \(\beta\) are the roots of the equation \(3x^2-4x+1=0\) the equation whose roots are \(\dfrac{\alpha^2}{\beta}\),\(\dfrac{\beta^2}{\alpha}\) is:-

OpenStudy (jiteshmeghwal9):

@sauravshakya @ParthKohli @UnkleRhaukus plz help :)

OpenStudy (anonymous):

One way to do it is by finding alpha and Beta

Parth (parthkohli):

\[\alpha\beta = \dfrac{1}{3}\]\[\alpha + \beta = \dfrac{4}{3}\]

OpenStudy (anonymous):

And find: (x-alpha^2/Beta) (x-Beta^2/Alpha) =0

Parth (parthkohli):

Or solve the equation, then use the fact that \((x - s)(x - r)\) is the equation is \(s\) and \(r\) are the roots.

Parth (parthkohli):

@jiteshmeghwal9 Didn't you read the `\dfrac` tip? =/

OpenStudy (jiteshmeghwal9):

I did it as follows :- \[3x^2-3x-x+1=0\]\[3x(x-1)-1(x-1)=0\]\[(3x-1)(x-1)=0\]\[\alpha=1, \beta=\frac{1}{3}\]\[\frac{\alpha^2}{\beta}=3\]\[{\beta^2 \over \alpha}=\frac{1}{9}\]\[3x^2+(3+\frac{1}{9})x+3.\frac{1}{9}\]

OpenStudy (jiteshmeghwal9):

yaa..i read it @ParthKohli

Parth (parthkohli):

It should be \(-\left(3 + \dfrac{1}{9}\right)\)

Parth (parthkohli):

Everything else is fine =)

OpenStudy (jiteshmeghwal9):

How did u gt a+b=4/3 ab=1/3 ? Is there any other method to find the roots ?

OpenStudy (anonymous):

if you observe the roots, the product of the new roots id \[(\alpha^{2}/\beta)*(\beta^{2}/\alpha)=\alpha \beta\] so the product of the roots doesnt change now find the some of the roots i.e. \[(\alpha^{3}+\beta^{3})/(\alpha \beta)=(\alpha+\beta)((\alpha+\beta)^{2}-3\alpha \beta)/(\alpha \beta)\] with this you can find the sum of the roots the quadratic equation is given by, \[x^{2}-(\alpha ^{1}+\beta^{1})x+\alpha^{1}\beta^{1}\] where alpha1 and beta1 are the new roots this is a generalized solution.

Parth (parthkohli):

I got that using Vieta's Formula. @jiteshmeghwal9 :P

OpenStudy (jiteshmeghwal9):

\[(x-3)(x-\frac{1}{9})=x^2-x/9-3x+1/3\] That's all i gt by putting the roots in the equation given by @sauravshakya but still i can't find the correct answer :(

Parth (parthkohli):

That's it!

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{green}{\text{Ans.-}3x^2-32x+3=0}}\]

OpenStudy (jiteshmeghwal9):

it's 3x^2-32x+3=0

OpenStudy (anonymous):

then the answer given must be wrong :)

Parth (parthkohli):

Yes, it's wrong.

OpenStudy (jiteshmeghwal9):

so what's the answer ?

OpenStudy (anonymous):

Did u get alpha =1 and Beta =1/3

OpenStudy (jiteshmeghwal9):

yup

OpenStudy (jiteshmeghwal9):

Look @ my first reply

Parth (parthkohli):

\[\left(x - \dfrac{1}{9}\right)\left(x -3\right)\]

OpenStudy (jiteshmeghwal9):

U mean the answer that i gt from saurav's equation is the answer

OpenStudy (jiteshmeghwal9):

@ParthKohli

OpenStudy (anonymous):

the process that i quoted earlier is also on of the standard processes. you can follow it. more over it is quite obvious that the product of the roots in the new equation must be the same as the earlier equation at 1/3. but the answers has a product of roots 1.

Parth (parthkohli):

Yes. You can get the answer from Vieta's Formula too.

OpenStudy (jiteshmeghwal9):

how ?

Parth (parthkohli):

The new roots are \(\dfrac{1}{9}\) and \(3\). So the prod. is \(\dfrac{1}{3}\) and the sum is \(\dfrac{28}{9}\). Now use Vieta's Formula!

OpenStudy (jiteshmeghwal9):

Thaanx... i gt it :)

Parth (parthkohli):

\[x^2 - \dfrac{28}{9}x + \dfrac{1}{3} =0\]You can multiply both sides by \(9\) to get a better equation.

OpenStudy (jiteshmeghwal9):

x^2-28x+3=0

Parth (parthkohli):

Actually \(9x^2\) :P

OpenStudy (jiteshmeghwal9):

right, LOL ;)